# Beamforming Design for Intelligent Reflecting Surface Aided Near-Field THz Communications

### v1

@@ IV Numerical Results
In this section, numerical results are presented to illustrate the effectiveness of the proposed DLDD IRS beamforming scheme. The center of the IRS is located at $(0,0,0)$ m of a three-dimensional coordinate system. The BS and the user are located at $(0,1.5,-1.5)$ m and $(2,-4,-2)$, respectively. It is assumed the IRS elements are separated by half of the wavelength, i.e., $d=\frac{\lambda_{c}}{2}$. Without otherwise specified, the parameters are set as: $N=100\times 100$, $K=10\times 10$, $f_{c}=300$, $B=30$, and $M=128$. For comparison, the optimal frequency-dependent IRS beamforming serves as the performance upper bound, where each element is equipped with a TD module [8], and the conventional IRS beamforming without TD modules is adopted as benchmark.
@@
The generated beams at the center frequency, $f_{c}$, and the edge subcarriers, $f_{1}$ and $f_{M}$, by adopting the proposed DLDD IRS beamforming scheme are presented in Fig. 2LABEL:sub@proposed_beamsplit. By comparing Fig. 2LABEL:sub@beamsplit and LABEL:sub@proposed_beamsplit, it is observed that the proposed scheme can focus the generated beams at different frequencies to the target location. The normalized array gains at different subcarriers are further illustrated in Fig. 4LABEL:sub@array_gain. It can be seen that the near-field beam split effect of IRS is suppressed by our proposed scheme. Specifically, in the wideband system with $B=30$ GHz, the normalized array gain of conventional narrowband IRS beamforming suffers over $90\%$ loss at the edge subcarriers, as compared to the narrowband case with $B=0.3$ GHz. By performing the proposed DLDD IRS beamforming scheme, the array gain loss can be effectively reduced to about $2\%$. Put differently, the proposed scheme can achieve up to $98\%$ near-optimal array gain across the entire bandwidth. Therefore, it can effectively serve as a wideband IRS beamforming design.
Fig. 4LABEL:sub@arraygain_vs_TTD shows the normalized edge subcarrier array gain versus the number of TD modules, $K_{t}$. We can see that even a small number of TD modules, e.g., $K_{t}=24$, readily achieves a good performance at edge subcarrier. Therefore, different strategies can be employed under different scenarios. For example, in the case of beam sweep, the angle spread is advantageous as it accelerates the process, whereas the TD modules may not be favorable. However, during data transmission in a large IRS-aided wideband communication system, the angle spread caused by the beam split is undesirable. This is because it significantly reduces the array gain at the edge frequencies compared to the center frequency. In this case, employing TD module is a favorable choice, and the proposed method can effectively reduce the required number of TD modules, which remarkably reduces the hardware cost and power consumption of the system.
Fig. 4LABEL:sub@arraygain_vs_tmax illustrates the normalized edge subcarrier array gain versus the required delay range, $t_{\text{req}}$. It is observed that the normalized edge subcarrier array gains of both schemes increase as $t_{\text{req}}$ scales up, as expected. However, the edge subcarrier array gain loss of the optimal solution proposed in [8] is large until $t_{\text{max}}$ is sufficiently large as mentioned in Section III, which is hard to realize in practice. In contrast, the proposed scheme achieves a normalized edge subcarrier array gain of $0.95$ for $t_{\text{req}}=9$ ps. This significant improvement comes from the DLDD IRS beamforming design, as it does not require the delay range to increase with distance.
In Fig. 4LABEL:sub@rate_pow, the achievable rate per subcarrier versus BS transmit power, $P_{\text{BS}}$, is provided. It is observed that the proposed scheme increases the achievable rate by about $110\%$ as compared to the narrowband design. Moreover, it approaches the optimal performance with a performance loss of less than $2\%$. Such slight performance loss arises from the design of the first layer TD network in eq. (18) as the phase difference between adjacent sub-surface along $y$-axis varies with different values of $k_{z}$. Nonetheless, the proposed scheme still achieves a near-optimal performance, which again reveals its superiority in alleviating the near-field beam split at the IRS.
@@
(c) Normalized edge subcarrier array gain versus $t_{\text{req}}$.
(d) Achievable rate per subcarrier versus $P_{\text{BS}}$.
@@
V Conclusion
In this letter, we proposed a novel wideband beamforming architecture that addresses the near-field beam spit effect for the IRS-aided THz communication system. Initially, the near-field beam split effect was analyzed, which revealed that the beams at different frequencies point towards different locations. To address this issue, we extended the piece-wise far-field model to the UPA-structured IRS, based on which, a DLDD IRS beamforming architecture was proposed. Simulation results demonstrated that the proposed DLDD IRS beamforming scheme can mitigate the near-field beam split effect and achieve a near-optimal performance with lower hardware cost. Notably, it showed that the maximum delay range required for TD modules is significantly reduced compared to the state-of-art delay-based architecture.

### v2

@@ V Numerical Results
In this section, numerical results are presented to illustrate the effectiveness of the proposed DLDD IRS beamforming and sub-optimal joint beamforming schemes. The center of the IRS is located at $(0,0,0)$ m of a three-dimensional coordinate system. The BS and the user are located at $(0,2,-1)$ m and $(2,-4,-2)$, respectively. It is assumed the IRS elements are separated by half of the wavelength, i.e., $d=\frac{\lambda_{c}}{2}$. Without otherwise specified, the parameters are set as: $N_{t}=32$, $K_{t}=4$, $N_{y}=N_{z}=80$, $K_{y}=K_{z}=10$, $f_{c}=300$ GHz, $B=30$ GHz, $M=128$, $P_{m}=20$ dBm, and $\sigma_{m}=-120$ dBm.
@@
To validate the effectiveness of the proposed DLDD IRS beamforming scheme, we first considered the single-antenna BS scenario, where $N_{t}=1$. The generated beams at the center frequency, $f_{c}$, and the edge subcarriers, $f_{1}$ and $f_{M}$, by adopting the proposed DLDD IRS beamforming scheme are presented in Fig. 2LABEL:sub@proposed_beamsplit. By comparing Fig. 2LABEL:sub@beamsplit and LABEL:sub@proposed_beamsplit, it is observed that the proposed scheme can focus the generated beams at different frequencies to the target location. The normalized array gains at different subcarriers are further illustrated in Fig. 4LABEL:sub@array_gain. It can be seen that the near-field beam split effect of IRS is suppressed by our proposed scheme. Specifically, in the wideband system with $B=30$ GHz, the normalized array gain of conventional narrowband IRS beamforming suffers nearly $100\%$ loss at the edge subcarriers, as compared to the narrowband case with $B=0.3$ GHz. By performing the proposed IRS beamforming scheme, the array gain loss can be significantly reduced to about $8\%$. Put differently, the proposed scheme can achieve up to $92\%$ near-optimal array gain across the entire bandwidth. Therefore, it effectively serve as a wideband IRS beamforming design.
Fig. 4LABEL:sub@arraygain_vs_TTD shows the normalized edge subcarrier array gain versus the number of TD modules, $K_{T}$. As expected, it increases with larger $K_{T}$. We can see that even a small number of TD modules, e.g., $K_{T}=63$, readily achieves a near-optimal performance. It is observed that the proposed scheme achieves the same performance as the sub-connected scheme in [6] with one fewer TD module, validating the calculation in Section III-C, where the proposed scheme requires $K\!-\!1$ TD modules, while the sub-connected scheme requires $K$. Although this difference is not significant, the reduced requirement on the range of the TD modules makes the proposed scheme more practical. This is illustrated in Fig. 4LABEL:sub@arraygain_vs_tmax, the proposed scheme achieves a normalized edge subcarrier array gain of $0.92$ with a required delay range of only $12$ ps, whereas both the optimal and sub-connected schemes rely on large-range TD modules, as discussed in Section III.
Then, the rate performance in multi-antenna BS scenario is provided, where the double-stage beam split effect is considered. Specifically, the average achievable rate per subcarrier based on our proposed joint beamforming scheme versus the system bandwidth, $B$, is plotted in Fig. 4LABEL:sub@rate_pow. It is revealed that the achievable rate of the conventional narrowband beamforming decreases rapidly as the bandwidth increases since it cannot address the double beam split effect, whereas the proposed scheme achieves enhanced rate performance. This is because the proposed joint beamforming can compensate the array gain loss caused by the double beam split effect even with large bandwidth. The impact of low-resolution IRS phase shifts is also provided by quantizing the obtained continuous phase shifts. It can be observed that the narrowband beamforming scheme suffers more performance loss from the discrete phase quantization than the proposed scheme, which implies the robustness of the latter against the practical hardware limitation. In fact, the $b=2$-bit phase shifters is an efficient choice in practical design, as its performance closely approaches that of continuous phase shifters, and it enjoys a lower complexity and hardware cost.
@@
(c) Normalized edge subcarrier array gain versus maximum delay range.
(d) Average achievable rate per subcarrier versus $B$.
@@
VI Conclusion
In this paper, we proposed a novel DLDD architecture that addresses the near-field beam spit effect for the IRS-aided THz communication system. Initially, the near-field beam split effect at the IRS was analyzed, which revealed that the beams at different frequencies point towards different locations. To address this issue, we proposed a DLDD beamforming architecture by extending the piecewise far-field model to the UPA-structured IRS. Then, we also proposed a sub-optimal joint beamforming scheme to address the double-stage beam split at the BS and the IRS. Simulation results demonstrated that the proposed DLDD IRS beamforming scheme can mitigate the near-field beam split effect and achieve a near-optimal performance with lower hardware cost. Besides, the proposed joint beamforming scheme effectively enhanced the rate performance with strong robustness across a wide bandwidth range.